# Syntax-Guided Procedural Synthesis of Molecules

### v1

@@ Appendix B Syntax Tree Recognition
In this section, we answer key questions like: (1) How does the relationship change with the addition of $\mathcal{T}$? (2) How strong is the correlation between $\mathcal{X}$ and $\mathcal{T}$? (3) How justified are the most confident predictions made by $Q_{\Theta}$? We investigate the relationship between $\mathcal{M}$ and $\mathcal{T}$. We seek to understand the extent to which the $Q\colon\mathcal{M}\rightarrow\mathcal{T}$ function is well-defined. The first part is quantitative analysis, and the second part is a qualitative study.
B.1 t-SNE and MDS Plots
@@
We use the t-distributed stochastic neighbor embedding (t-SNE) on the final layer hidden representations of our MLP $Q_{\Theta}$ to visualize how our recognition model discriminates between molecules of different syntax tree classes. From 6, we see the MLP is able to discriminate amongst the top 3 or 4 most popular skeleton classes, visually partitioning the representation space. However, beyond that the representations on the validation set begin to coalesce, i.e., the model begins overfitting.
@@
Since gradient descent is stochastic, we also use multi-dimensional scaling (MDS) using the Morgan Fingerprint Manhattan distance on a subset of our dataset to visualize the relative positioning between molecules of different syntax tree classes (sorted based on popularity). From the plots in 8, we observe some interesting trends:
@@
  • The classes do not form disjoint partitions in space. As the ranked popularity increases, the points tend to disperse outwards. There are exception classes, e.g., the yellow set of points in 7(b) that cluster in the center.
@@
Figure 8: We visualize the structure-property relationship as a scatterplot of 2D structures vs property values. (Top) Structure is $\mathcal{T}\times\mathcal{X}$. We use MDS with the dissimilarity $\text{dist}((T_{1},X_{1}),(T_{2},X_{2}))=\text{Tree-Edit-Distance}(T_{1},T_{2})+\text{Manhattan}(X_{1},X_{2})$. (Bottom) Structure is only $\mathcal{X}$.
We also use MDS to investigate the structure-property relationship to understand the joint effect $\mathcal{T}$ and $\mathcal{X}$ has on different properties of interest. As shown in 8, we see overall, the functional landscape varies significantly from property to property, but the general trend is that decoupling $\mathcal{T}$ from $\mathcal{X}$ does not change the structure-property relationship much. Whereas analog generation requires a more granular understanding of the synergy between $\mathcal{X}$ and $\mathcal{T}$, molecular optimization does not. Instead, the evolutionary strategy should be kept fairly consistent between the original design space $(\mathcal{X})$ and $(\mathcal{X}\times\mathcal{T})$. However, the top row exhibits lower entropy, with the empirical distribution looking “less Gaussian". To capture this nuance, the evolutionary algorithm should combine both global and local optimization steps. We meet this observation with a bilevel optimization strategy that combines Semantic Crossover with Syntactic Mutation.
B.2 Expert Case Study
In this section, we enter the perspective of the recognition model learning the mapping from molecules to syntax tree skeletons. The core difference between this exercise and a common organic chemistry exam question [19] is the option to abstract out the specific chemistry. Since the syntax only determines the skeletal nature of the molecule, the specific low-level dynamics don’t matter. As long as the model can pick up on skeletal similarities between molecules, it will be confident in its prediction. We did the following exercise to understand if cases where the recognition model is most confident on unseen molecules can be attributed to training examples. We took the following steps:
@@ B.2 Expert Case Study
Shown in Figures 9 and 10 is the output of these steps for a common skeleton class which requires two reaction steps.
@@
In Figure 9, we see that the query molecule’s nearest neighbor is an output from the same program but different building blocks. Both feature the same core fused ring system involving a nitrogen. Given that the model has seen 9(b) (and other similar instances), it should associate this core feature with a ring formation reaction step. Taking a step deeper, the respective precursors also share the commonality of having an amide linkage in the middle. Amides are key structural elements that the recognition model can identify. Both precursors underwent the same amide linkage formation step, despite the building blocks being different. Thus, the model’s high confidence on the query molecule can be attributed directly to 9(b).
In Figure 10, there is more “depth" to the matter. We see a skeletal similarity across all three molecules: a nitrogen in the center with three substituents. Although it’s noteworthy that the nitrogen participates in a sulfonamide group in all three cases, using this fact to inform the syntax tree would be a mistake. This is because in 10(a) and 10(b), the sulfonamide group is the result of an explicit sulfonamide formation reaction, where a sulfonyl chloride reacts with an amine. However, in 10(c) the sulfonamide group is already present in a building block. Thus, we see where the recognition model taking as input the circular fingerprint of this molecule could overfit. Nonetheless, the nitrogen with three substituents necessitates at least one reaction is required. The necessity for a second reaction can be attributed to the ether linkage present in both 10(a) and 10(c). The recognizer would be able to justify an additional reaction after it has seen the bicyclic ring structure joined with the sulfonamide group sufficiently many times before. In summary, the model will often be presented multiple complex motifs, but only a subset of them may be responsible for reaction steps. The exact number of reactions needed can only be determined via actually doing the search, but high-level indicators (such as the nitrogen with three substituents) allow the recognition model to abstract out the semantic details and construe a “first guess" of what the syntax tree is.
Appendix C Connection with Program Synthesis
@@
Program synthesis is often formulated as deducing a constructive proof to the statement: for all inputs, there exists an output such that $\phi$ holds. The constructive proof itself is then the program. At the low-level, program synthesis methods repeatedly calls a SAT solver with the logical formula $\neg\phi$. If UNSAT is returned, this means $f$ is valid. Syntax-guided synthesis [1, 53] (SyGuS) is a framework for meeting the correctness specification with a syntactic template. Syntactic templates explicitly constrains $G_{L}$, significantly reducing the number of implementations $f$ can take on. Sketching is an example application where programmers can sketch the skeletal outline of a program for synthesizers to fill in the rest [57]. More directly related to our problem’s formulation is inductive synthesis, which seeks to generate $f$ to match input/output examples. The problem of synthesis planning for a molecule $M$ is a special case of the programming-by-example paradigm, where we seek to synthesize a program consistent with a single input/output pair: ($\{B\}$, $M$). Inductive synthesis search algorithms have been developed to search through the combinatorial space of derivations of $G_{L}$. In particular, stochastic inductive synthesis use techniques like MCMC to tackle complex synthesis problems where enumerative methods do not scale to. MCMC has been used to optimize for the opcodes in a program [53] or for the Abstract Syntax Tree (AST) directly [1]. In our case, the space of possible program semantics is so large that we decouple the syntax from the semantics, performing stochastic synthesis over only the syntax trees. We also borrow ideas from functional program synthesis, where top-down strategies are preferred over bottom-up ones to better leverage the connection between a high-level specification and a concrete implementation [49]. Similar to how top-down synthesis enables aggressive pruning of the search space via type checking, retrosynthesis algorithms leverages the target molecule $M$ to prune the search space via template compatability checks.
Appendix D Derivation of Grammar
@@
  • How can when pick the initial production rule when the number of syntax trees grow large? We use an iterative refinement strategy, governed by a Markov Chain Process over the space of syntax trees. The simulation is initialized at the structure predicted from our recognition model B.
@@
These two questions are addressed by the design choices in 3.3.
Appendix E Policy Network
E.1 Featurization
Recall $\Phi\colon T^{\prime}\rightarrow\mathbb{R}^{|T^{\prime}|\times 91}$ and $\Omega\colon T^{\prime}\rightarrow\mathbb{R}^{|T^{\prime}|\times 256}$. Our dataset $D_{\text{pretrain}}$ consists of instances $(M^{(i)},{T^{\prime}}^{(i)},X^{(i)},y^{(i)})$. We adopt Morgan Fingerprints with radius 2 as our encoder (FP). Then, we featurize each instance as:
$X^{(i)}_{n}$ $\coloneqq[\text{FP}_{2048}(M^{(i)});\text{BB}(n);\text{RXN}(n)],\qquad$

### v2

@@ Appendix C Syntax Tree Recognition
In this section, we answer key questions like: (1) How does the relationship change with the addition of ${\mathcal{T}}$? (2) How strong is the correlation between ${\mathcal{X}}$ and ${\mathcal{T}}$? (3) How justified are the most confident predictions made by $\tau{}{}$? We investigate the relationship between $\mathcal{M}$ and ${\mathcal{T}}$. We seek to understand the extent to which the true mapping $\mathcal{M}\rightarrow{\mathcal{T}}$ is well-defined. The first part is quantitative analysis, and the second part is a qualitative study.
C.1 t-SNE and MDS Plots
@@
We use the t-distributed stochastic neighbor embedding (t-SNE) on the final layer hidden representations of our MLP $\tau{}{}$ to visualize how our recognition model discriminates between molecules of different syntax tree classes. From Figure 7, we see the MLP is able to discriminate amongst the top 3 or 4 most popular skeleton classes, visually partitioning the representation space. However, beyond that the representations on the validation set begin to coalesce, i.e., the model begins overfitting.
@@
Since gradient descent is stochastic, we also use multi-dimensional scaling (MDS) using the Morgan Fingerprint Manhattan distance on a subset of our dataset to visualize the relative positioning between molecules of different syntax tree classes (sorted based on popularity). From the plots in Figure 9, we observe some interesting trends:
@@
  • The classes do not form disjoint partitions in space. As the ranked popularity increases, the points tend to disperse outwards. There are exception classes, e.g., the yellow set of points in Figure 8(b) that cluster in the center.
@@
Figure 9: We visualize the structure-property relationship as a scatterplot of 2D structures vs property values. (Top) Structure is ${\mathcal{X}}\times{\mathcal{T}}$. We use MDS with the dissimilarity $d_{{\mathcal{X}}\times{\mathcal{T}}}(({\bm{x}}_{1},T_{1}),({\bm{x}}_{2},T_{2}))=||{\bm{x}}_{1}-{\bm{x}}_{2}||_{1}+\text{Tree-Edit-Distance}(T_{1},T_{2})$. (Bottom) Structure is only ${\mathcal{X}}$.
We also use MDS to investigate the structure-property relationship to understand the joint effect ${\mathcal{T}}$ and ${\mathcal{X}}$ has on different properties of interest. As shown in Figure 9, we see overall, the functional landscape varies significantly from property to property, but the general trend is that decoupling ${\mathcal{T}}$ from ${\mathcal{X}}$ does not change the structure-property relationship much. Whereas analog generation requires a more granular understanding of the synergy between ${\mathcal{X}}$ and ${\mathcal{T}}$, molecular optimization does not. Instead, the evolutionary strategy should be kept fairly consistent between the original design space $({\mathcal{X}})$ and $({\mathcal{X}}\times{\mathcal{T}})$. However, the top row exhibits lower entropy, with the empirical distribution looking “less Gaussian". To capture this nuance, the evolutionary algorithm should combine both global and local optimization steps. We meet this observation with a bilevel optimization strategy that combines semantic crossover with syntactic mutation.
C.2 Expert Case Study
In this section, we enter the perspective of the recognition model learning the mapping from molecules to syntax tree skeletons. The core difference between this exercise and a common organic chemistry exam question (Flynn, 2014) is the option to abstract out the specific chemistry. Since the syntax only determines the skeletal nature of the molecule, the specific low-level dynamics don’t matter. As long as the model can pick up on skeletal similarities between molecules, it will be confident in its prediction. We did the following exercise to understand if cases where the recognition model is most confident on unseen molecules can be attributed to training examples. We took the following steps:
@@ C.2 Expert Case Study
Shown in Figures Figure 10 and Figure 11 is the output of these steps for a common skeleton class which requires two reaction steps.
@@
In Figure 10, we see that the query molecule’s nearest neighbor is an output from the same program but different building blocks. Both feature the same core fused ring system involving a nitrogen. Given that the model has seen Figure 10(b) (and other similar instances), it should associate this core feature with a ring formation reaction step. Taking a step deeper, the respective precursors also share the commonality of having an amide linkage in the middle. Amides are key structural elements that the recognition model can identify. Both precursors underwent the same amide linkage formation step, despite the building blocks being different. Thus, the model’s high confidence on the query molecule can be attributed directly to Figure 10(b).
In Figure 11, there is more “depth" to the matter. We see a skeletal similarity across all three molecules: a nitrogen in the center with three substituents. Although it’s noteworthy that the nitrogen participates in a sulfonamide group in all three cases, using this fact to inform the syntax tree would be a mistake. This is because in Figure 11(a) and Figure 11(b), the sulfonamide group is the result of an explicit sulfonamide formation reaction, where a sulfonyl chloride reacts with an amine. However, in Figure 11(c) the sulfonamide group is already present in a building block. Thus, we see where the recognition model taking as input the circular fingerprint of this molecule could overfit. Nonetheless, the nitrogen with three substituents necessitates at least one reaction is required. The necessity for a second reaction can be attributed to the ether linkage present in both Figure 11(a) and Figure 11(c). The recognizer would be able to justify an additional reaction after it has seen the bicyclic ring structure joined with the sulfonamide group sufficiently many times before. In summary, the model will often be presented multiple complex motifs, but only a subset of them may be responsible for reaction steps. The exact number of reactions needed can only be determined via actually doing the search, but high-level indicators (such as the nitrogen with three substituents) allow the recognition model to abstract out the semantic details and construe a “first guess" of what the syntax tree is.
Appendix D Expanded Related Works
D.1 Background on Program Synthesis
@@
Program synthesis is often formulated as deducing a constructive proof to the statement: for all inputs, there exists an output such that $\phi$ holds. The constructive proof itself is then the program. At the low-level, program synthesis methods repeatedly calls a SAT solver with the logical formula $\neg\phi$. If UNSAT is returned, this means $f$ is valid. Syntax-guided synthesis (Alur et al., 2013; Schkufza et al., 2013) (SyGuS) is a framework for meeting the correctness specification with a syntactic template. Syntactic templates explicitly constrains $G_{L}$, significantly reducing the number of implementations $f$ can take on. Sketching is an example application where programmers can sketch the skeletal outline of a program for synthesizers to fill in the rest (Solar-Lezama et al., 2005). More directly related to our problem’s formulation is inductive synthesis, which seeks to generate $f$ to match input/output examples. The problem of synthesis planning for a molecule $M$ is a special case of the programming-by-example paradigm, where we seek to synthesize a program consistent with a single input/output pair: ($\{B\}$, $M$). Inductive synthesis search algorithms have been developed to search through the combinatorial space of derivations of $G_{L}$. In particular, stochastic inductive synthesis use techniques like MCMC to tackle complex synthesis problems where enumerative methods do not scale to. MCMC has been used to optimize for the opcodes in a program (Schkufza et al., 2013) or for the abstract syntax tree directly (Alur et al., 2013). In our case, the space of possible program semantics is so large that we decouple the syntax from the semantics, performing stochastic synthesis over only the syntax trees. We also borrow ideas from functional program synthesis, where top-down strategies are preferred over bottom-up ones to better leverage the connection between a high-level specification and a concrete implementation (Polikarpova et al., 2016). Similar to how top-down synthesis enables aggressive pruning of the search space via type checking, retrosynthesis algorithms leverages the target molecule $M$ to prune the search space via template compatability checks.
D.2 Execution-guided Program Synthesis
We would like to note the distinction between our program synthesis formulation and other formulations. Retrosynthesis is essentially already guided by the execution state at every step. Each expansion in the search tree executes a deterministic reaction template to obtain the new intermediate molecule. Planners based on single-step models (Chen et al., 2020), for example, assume the Markov Property by training models to directly predict a template given only the intermediate (Torren-Peraire et al., 2024; Tu et al., 2022). In program synthesis, meanwhile, the state space is a set of partial programs with actions corresponding to growing the program. The execution of the program (or verification against the specification) does not happen until a complete program is obtained. In recent years, neural program synthesis methods found using auxiliary information in the form of the execution state of a program can help indirectly inform the search (Bunel et al., 2018; Chen et al., 2018; Ellis et al., 2019) since it gives a sense on what the program can compute so far. This insight does not apply to retrosynthesis, since retrosynthesis already executes on the fly. It also does not apply for the methods introduced in Section 2.2 that construct a synthetic tree in a bottom-up manner, for the same reason (the only difference is they use forward reaction templates, with a much smaller set of robust reaction templates) to obtain the execution state each step. However, as described in Section 3.3.1, our approach combines the computational advantages of restricting to a small set of forward reaction templates with the inductive bias of retrosynthetic analysis. Our policy is to predict forward reaction templates in a top-down manner. This formulation is common in top-down program synthesis, where an action corresponds to selecting a hole in the program. Similarly, our execution of the program does not happen until the tree is filled in. However, we leverage the insight that the execution state helps in an innovative way, as discussed in F.3.
D.3 Inspirations from Retrosynthesis and Alternate Formulations
We begin by elaborating the distinctions between retrosynthesis methods and methods for synthesizable molecular design. Then, we identify a few recent works from retrosynthesis that can inspire cross-pollination of ideas. Finally, we end with alternate formulations of the problem that are also valuable to consider for future cross-examination.
Intended Use Case
Retrosynthesis aims to find a synthetic route for a given target molecule, without reference to how the target molecule is obtained or further optimizations on the target molecule. The target molecule is a compound that may serve any application or use case that we will not get into here, but importantly it is the problem to solve. We refer readers to Gao & Coley (2020) for further descriptions.
Synthesizable molecular design aims to be a standalone molecular optimization workflow that explicitly constrains the design space to be synthetically accessible building blocks and reactions Vinkers et al. (2003). This is often coupled with property oracles that evaluate the designs, which guides the optimization towards parts of the design space with higher fitness Gao et al. (2022).
MDP Formulation
Retrosynthesis can be formulated as a tree-shaped MDP, where each state is a molecule (initial state being the target, terminal states being building blocks) and each action is a reversed (“retro") reaction. The tree shape of the MDP is due to the fact the retro reaction (action) produces a set of reactants (states) Liu et al. (2023). Retrosynthetic planners often tackle the MDP by combining a single-step model (predicting retro reactions) and a multi-step planner (e.g. A* search Liu et al. (2018), MCTS Segler & Waller (2017), depth-first search Kishimoto et al. (2019)). A solution to the MDP is a tree of actions, i.e. a synthetic tree, where all sequences of actions in this tree lead to terminal states.
Synthesizable molecular design feature a broader set of methods but can be defined as a discrete optimization problem over synthetic trees directly: $\operatorname*{arg\,max}_{x}f(F(x))$ where $x$ is a synthetic tree, $F$ the root molecule of $x$ and $f$ is the fitness function. As the design space is intractably large, prior approaches discussed in 2.3 formulate the problem as a serial MDP, where each state is a synthetic tree and each action an edit operation (add, merge, etc.) to the synthetic tree. Though simple, we argue such a formulation is ill-advised, for reasons we discussed and demonstrated in the main text. We use an alternate formulation, inspired by program synthesis, that considers each state as a partially sketched program and each action as completing a hole of the program.
Learning Goal
Retrosynthesis methods aim to learn a policy $\pi(a|s)$, where $s$ is a molecule and $a$ a retro reaction, which can be template-based or template-free (we won’t go into that here). Traditional works learn the policy from public datasets of synthetic routes (e.g. USPTO), but recent works have explored novel strategies for learning $\pi$ by combining offline and online training. The offline training is usually done on a reaction dataset to initialize a policy network and/or reaction model. The online training iteratively adapts the policy network by acquiring more data using a planning algorithm, possibly guided by the current policy network. More specifically, Guo et al. (2024) uses MCTS to acquire data, inferring policy and value targets based on the node visit counts. Kim et al. (2021) uses a self-improvement strategy, reminiscent of AlphaGo Zero Silver et al. (2017), that trains independent reference forward and backward reaction models to control the quality and diversity of new reaction pathways acquired by the planner. Liu et al. (2023) follows a similar strategy, but decouples the synthesizability and cost of the value function. They also architect a two-branch policy network that uses a trainable single-step network to optimize the probabilities over reactions from the frozen reference network to better model real-world synthesis considerations. Like Guo et al. (2024), they use MCTS guided by the current policy network to acquire more data, creating a synergistic feedback loop that results in a holistic, trained policy network which can be plugged into multi-step planners.
Synthesizable molecular design methods, meanwhile, are relatively more spread out, with research going into problem formulation and algorithmic frameworks for tackling this more open-ended problem. In SynthesisNet, the policy learning is done entirely offline (as described in Sections 3.3.1 and 4.1.1) to amortize for the cost of searching during the actual online phases (MCMC and GA), but above techniques from retrosynthesis can also facilitate self-improving the surrogate network. One potential idea for generating more experiences is to take partial program examples from the existing data, use guided planning to complete those examples, then retrain the surrogate network with the augmented dataset. We leave the details for exciting future extensions of our work.
Alternate Formulations
Towards synthesizable molecular design, alternate formulations from recent developments can also be considered. Decision Transformer Chen et al. (2021) is a recent work that re-imagines offline RL as a conditional sequence modeling task. Notably, the model conditions on reward-to-go and the history for generating the next action. The Transformer architecture enables long-term modeling of the environment’s dynamics, enabling credit assignment and relational modeling of its history. The offline setting tackled by Decision Transformer naturally aligns with our method, where we do (self-)supervised policy training from programs generated offline. Although the program structure carries important hierarchical information about the relations of building blocks and reactions to one another, it’s worth considering whether the synthesis tree construction serialization protocol used by prior works Bradshaw et al. (2020); Gao et al. (2021); Luo et al. (2024); Gao et al. (2024) can be used to formulate a conditional sequence modeling problem. GFlowNet Bengio et al. (2021; 2023) is also a recent work that formulates a flow network over a tree-based MDP, where the incoming and outgoing flow are proportional the probability of subsequent actions and learned using flow matching objectives. The goal of this model is to learn amortized samplers for reward functions, producing both diverse and high-quality samples constructed in a step-by-step manner following an MDP environment. The learning occurs on offline trajectories with observed rewards. The formulation using reward functions can be a direction of future work for our framework, which currently only considers rewards in the online phases. However, if we had considered rewards to be provided upfront along with the data generation procedure, we can adopt GFlowNet to amortize the expensive work done by MCMC, and directly sample programs. We leave this study for future works. We believe cross-examining alternate formulations and recent methodologies to be essential for finding future inspirations for extending the innovation horizon of methods used to tackle synthesizable molecular design.
Appendix E Derivation of Grammar
@@
  • How can when pick the initial production rule when the number of syntax trees grow large? We use an iterative refinement strategy, governed by a Markov Chain Process over the space of syntax trees. The simulation is initialized at the structure predicted from our recognition model Appendix C.
@@
These two questions are addressed by the design choices in Section 3.3.
Appendix F Policy Network
F.1 Featurization
Our dataset ${\mathcal{D}}$ comprises partial programs ${T\in\partial{\mathcal{P}}}$ producing molecules $M$.Then, we compute node features ${\bm{H}}$ and labels ${\bm{Y}}$ as:
${\bm{h}}_{n}\coloneqq[\text{FP}_{2048}(M),\text{BB}_{2048}(n),\text{RXN}(n)],\quad{\bm{y}}_{n}\coloneqq\begin{cases}\text{RXN}(n),&\text{if $i$ is a reaction node},\\ \text{BB}_{256}(n),&\text{otherwise},\end{cases}$
where $\text{FP}_{d}(\cdot)$ computes the $d$-bit radius 2 Morgan fingerprints, $\text{BB}_{d}(n)=\text{FP}_{d}(n_{\text{SMILES}})$ if $n$ is attributed with a building block from ${\mathcal{B}}$ or $\textbf{0}_{d}$ otherwise and $\text{RXN}(n)=\text{one\_hot}_{91}(n_{\text{\text{RXN}\_ID}})$ if $n$ is attributed with a reaction from ${\mathcal{R}}$ or $\textbf{0}_{91}$ otherwise.
If $\mathcal{N}(T)$ and $\mathcal{E}(T)$ denote the node and edge set of $T\in\partial{\mathcal{P}}$, then we define, for convenience:
$\text{RXN}({T})$ :=$:=\{r\in\mathcal{N}(T)\mid\exists c,p\in\mathcal{N}(T)\text{ s.t.}(r,c)\in\mathcal{E}(T)\cap(p,r)\in\mathcal{E}(T)\}$ (1)
$\text{BB}({T})$ =$=\{b\in\mathcal{N}(T)\mid\nexists c\text{ s.t. }(b,c)\in\mathcal{E}(T)\}.\}$ (2)
F.2 Loss Function
Let the superscript $(i)$ indicate the $i$-th sample in the dataset. The loss function is:
$\mathcal{L}_{{\mathcal{D}}}(\Phi)$ $\coloneqq\frac{1}{|{\mathcal{D}}|}\sum_{i=1}^{|{\mathcal{D}}|}\sum_{n\in\text{RXN}(T^{(i)})}\text{CE}(\pi_{\mathcal{R}}(T^{(i)},{\bm{H}}^{(i)})_{n},{\bm{y}}^{(i)}_{n}),$
$\mathcal{L}_{{\mathcal{D}}}(\Omega)$ $\coloneqq\frac{1}{|{\mathcal{D}}|}\sum_{i=1}^{|{\mathcal{D}}|}\sum_{n\in\text{BB}(T^{(i)})}\text{MSE}(\pi_{\mathcal{B}}(T^{(i)},{\bm{H}}^{(i)})_{n},{\bm{y}}^{(i)}_{n}).$
CE and MSE denote the standard cross entropy loss and mean squared error loss, respectively. For our evaluation metric, we consider accuracy, where the output of $\pi_{{\mathcal{B}}}$ is interpreted as the nearest building block with respect to cosine distance.
F.3 Auxiliary Training Task
In Section 3.3.1, we defined the representation $T$ to be the parse tree of a partial program. However, we omitted an extra step that was used to preprocess $T$ for training. The motivation for this extra step is discussed deeply in D.2. We add an additional step when preprocessing ${\mathcal{D}}$: For each $T$ in ${\mathcal{D}}$, for each node $r$ corresponding to a reaction, we add a new node $o_{r}$ corresponding to the intermediate outcome of the reaction. If $\text{RXN}(T)$ is the reaction nodes of $T$, we can construct $T^{\prime}$ from $T$ as follows:
$\mathcal{N}(T^{\prime})$ $\leftarrow\mathcal{N}(T)\cup\text{RXN}(T)$ (3)
$\mathcal{E}(T^{\prime})$ $\leftarrow\mathcal{E}(T)\cup\{(\text{parent}(r),o_{r}),(o_{r},r)\forall r\in\text{RXN}(T)\}$ (4)
Figure 12: Examples of $T^{\prime}$ where prediction targets are the frontier reactions (yellow circles), frontier building blocks (numbered yellow squares) and auxiliary intermediates (un-numbered yellow squares).
Lastly, we attribute each $o_{r}$ with the intermediate obtained from the original synthetic tree, i.e. executing the output of the program rooted at $r$. We featurize $\{y_{o}:=\text{FP}_{256}({o}_{\text{SMILES}})\}$ and add them as additional prediction targets to ${\mathcal{D}}$. Examples of $T^{\prime}$ are given in Figure 12.
F.4 Ablation Study: Auxiliary Task
To understand whether the two key design choices for $\partial{\mathcal{P}}^{\prime}$ are justified, we did two ablations:
  1. 1.
    We use the original description of $\partial{\mathcal{P}}$ in Section 3.3.1, i.e. without the auxiliary task.
  2. 2.
    We use $\partial{\mathcal{P}}^{\prime}$, but without attributing the intermediate nodes (so the set of targets is the same as Ablation 1.)
(a) Examples from $\partial{\mathcal{P}}^{\prime}$
(b) NN accuracy loss over top example Figure 15(a)
(c) NN accuracy over bottom example Figure 15(a)
(d) NN accuracy over the validation set
As shown in Figure 13(d), using $\partial{\mathcal{P}}^{\prime}$ (Ours) achieves higher NN accuracy. This shows the benefit of learning the auxiliary training task. Meanwhile, ablating the auxiliary task (-aux) and ablating the intermediate node (-interm) does not have meaningful difference, indicating our architecture is robust to graph edits which are semantically equivalent. To understand the comparative advantage vs disadvantage of the auxiliary training task, consider the two examples in Figure 15(a). The first example is equivalent to learning a single-step backward reaction prediction on forward templates [^3] For some templates, the forward template is one-to-one. For others, applying the backward template results in an ill-defined precursor, due to the many-to-one characteristic of these templates.. Our model clearly benefits from the auxiliary training task, which provides additional examples for learning the backward reaction steps. However, our model fares worse on predicting the first reactant of the top reaction. This may be due to competing resources. Despite the task being the same (and the set of forward templates are fixed), the model has to allocate sufficient capacity for the auxiliary task, whose output domain is much higher dimensional than ${\mathcal{B}}$. Ensuring positive transfer from learning the auxiliary task is an interesting extension for future work.
F.5 Ablation Study: Linear Training
Table 6: We follow the same setup as Table 1, but retrain Ours using a dataset constructed with parameter $k=4$ (whereas we used $k=3$ in Table 1) to match the linear sampling strategy models, which we refer to as Ours:MC. We evaluate models trained using sampling constants 1 and 10, as described in 4.1.3.
Avg. Sim. $\uparrow$ SA $\downarrow$ Diversity $\uparrow$
Dataset Method RR $\uparrow$ Top-1 Top-3 Top-5 Top-1 Top-3 Top-5 Top-3 Top-5
Test Set Ours ($\tau$) 56% 0.827 0.633 0.555 3.100 3.019 2.918 0.543 0.628
Ours:MC1 ($\tau$) 36% 0.732 0.564 0.513 3.048 2.913 2.844 0.609 0.665
Ours:MC10 ($\tau$) 65% 0.869 0.658 0.609 3.163 3.000 2.928 0.558 0.610
ChEMBL Ours ($\tau$) 7.6% 0.531 0.443 0.396 2.544 2.510 2.460 0.675 0.727
Ours (MCMC) 9.2% 0.532 0.486 0.432 2.364 2.310 2.263 0.765 0.759
Ours:MC1 (MCMC) 2.0% 0.406 0.337 0.289 2.604 2.563 2.439 0.756 0.767
Ours:MC10 (MCMC) 8.5% 0.519 0.421 0.367 2.644 2.420 2.331 0.618 0.640
F.5.1 Results on Synthesizable Analog Generation
We study whether the efficiency of a sampling-based training strategy comes at a cost of performance. We make two observations from the results in Table 6.
Constant factor matters. The constant multiplier $C$ from $\mathcal{D}_{0}$ to $\mathcal{D}$ not only determines how much data the model sees each pass for the sake of efficiency. It is also be a parameter controlling the tradeoff between over-representing larger vs smaller templates. If it is larger than the largest number of masks (Table 5), the Linear strategy is essentially deactivated, since all masks will be used. At a lower value, only small programs with at most the number of masks as $C$ are fully represented. Medium-to-larger programs in $\mathcal{D}_{0}$ are under-represented, at the rate of the fraction of total masks that $C$ constitutes for its template class.
Performance boost in-distribution. We find that for $C=10$, the performance is better across reconstruction, similarity and diversity with comparable SA to the standard training. Meanwhile for $C=1$, the performance declines sharply. It is likely that standard training is overfitting to masks from larger programs, resulting in poorer generalization. Meanwhile, $C=10$ downsamples those programs, and its sampling can be viewed as data-level regularization against overfitting.
Slight performance drop out-of-distribution. We find both $C=1$ and $C=10$ underperform compared to standard training. For $C=10$, reconstruction and Top-1 similarity are actually comparable, but its similarity, SA and diversity are noticeably worse than standard training. Since ChEMBL feature predominantly unsynthesizable molecules, it is likely that the distribution of molecular fingerprints better reflect those outputs of the more complex programs in $\mathcal{D}_{0}$, which are downweighted by higher values of $C$.
F.5.2 Results on Synthesizable Molecular Design
Figure 14: We select the first Oracle from each Table in App. H to compare Ours with Ours (EP). Aside from the ablation networks, we use the same experimental settings as Table 2.